OpenStudy (anonymous):

g(x)=x^2+1 shifted two units to right

OpenStudy (whpalmer4):

Are you supposed to graph it, write the equation, what?

OpenStudy (anonymous):

write the equation

OpenStudy (rulnick):

To shift to the right, just replace x with x-2.

OpenStudy (anonymous):

x-2^2+1 ?

OpenStudy (rulnick):

g(x) = (x-2)^2 + 1

OpenStudy (whpalmer4):

No, you have to square the whole (x-2)

OpenStudy (rulnick):

simplified, it's x^2-4x+5

OpenStudy (anonymous):

okay thank you

OpenStudy (rulnick):

welcome

OpenStudy (anonymous):

to move to the left, would it be (x+2)^2+1

OpenStudy (anonymous):

?

OpenStudy (whpalmer4):

Yes. In general, to move right or left, you subtract or add from the argument to the function. Take y = x^2: at x = 0, y = 0^0 = 0. If you subtract 1 from x before squaring it, that means that the spot where y = 0 is moved by 1 unit along the x axis, and x = 1 is where the feature that happened at x=0 previously now occurs. Does that make sense?

OpenStudy (anonymous):

yes! that actually makes alot of sense. this isnt related, but could you explain to me how to tell if a graph is a function ?

OpenStudy (whpalmer4):

Yes, I can :-) If there is any spot on the graph where you can draw a vertical line and cross the curve twice, it is not a function. A function is defined as only having 1 value of y for any given value of x.

OpenStudy (anonymous):

so a parabola would be a function ?

OpenStudy (whpalmer4):

If it is a parabola of the sort that opens up or down, yes. One that opens right or left, no. y=x^2 is a function, x = y^2 is not.

OpenStudy (anonymous):

okay, that makes sense. thank you for explaining to me

OpenStudy (whpalmer4):

The inverse of a function is the function reflected about the line y = x, so an inverse of a function is not a function.

OpenStudy (rulnick):

That last statement is not true. If it were, there would be no one-to-one functions.

OpenStudy (whpalmer4):

give me an example of what you mean.

OpenStudy (whpalmer4):

sorry, you're right, my statement is overly broad there.

OpenStudy (rulnick):

Example: y=2x is a function (y is a function of x), so is the inverse x=y/2 (x is a function of y); or, if you prefer, reversing the variable mapping, y=x/2 (y is a function of x); either way.

OpenStudy (rulnick):

np, good thoughts nonetheless

OpenStudy (whpalmer4):

no inverse of a function where a horizontal line crosses the graph with multiple intersections...

OpenStudy (whpalmer4):

(essentially the inverse of the vertical line test)

OpenStudy (whpalmer4):

the inverse of a function where a horizontal line crosses the graph with multiple intersections is not a function...that's what I mean to say :-)

OpenStudy (anonymous):

i have another question, im sorry !

OpenStudy (whpalmer4):

answering your questions is why I'm here!

OpenStudy (anonymous):

g(x)=x^3 stretched to be twice as tall, im confused because its a cubic function

OpenStudy (whpalmer4):

Okay, I meant to touch on this too when we were talking about shifting right and left. If you want to shift up and down, you add or subtract to the result of the function. Agreed?

OpenStudy (anonymous):

yes!

OpenStudy (whpalmer4):

It's like printing the function on a transparency and sliding it up and down the y-axis.

OpenStudy (anonymous):

what does stretch mean ?

OpenStudy (whpalmer4):

now, stretching (or compressing) along the y-axis is similar to printing on transparent rubber...think of what happens to a balloon with printing on it when you blow it up.

OpenStudy (whpalmer4):

what would be the mathematical operation to do that, do you think?

OpenStudy (anonymous):

hmm ... muliplication. the print would get bigger

OpenStudy (whpalmer4):

yes!

OpenStudy (anonymous):

am i wrong ?/:

OpenStudy (anonymous):

aww yay

OpenStudy (whpalmer4):

and multiplication by a fraction would make it smaller.

OpenStudy (whpalmer4):

Now, here's a trickier one: how do you stretch or compress along the x-axis?

OpenStudy (anonymous):

uhhh .... divide ? assuming

OpenStudy (whpalmer4):

well, maybe. The subtle point here is that like we did to shift right and left, we change the argument to the function, instead of the result of the function...

OpenStudy (whpalmer4):

Think about that for a minute...

OpenStudy (whpalmer4):

okay, time's up! :-)

OpenStudy (anonymous):

wait - so x^3 stretched twice it's height would be, x^3(2)

OpenStudy (whpalmer4):

Say we have a sine wave...y = sin x:  Now, if we change that to be y = sin 2x, the curve wiggles twice as many times in a given amount of x-axis, right?